I came across this IQ question:
How do you proof that 10 = 9.999999999...?

Answer by venugopalramana(3286)

How do you proof that 10 = 9.999999999...?
LET ......X = 9.9999999999999.....
SO ....10X = 99.99999999999999.........SUBTRACTING
10X-X=90
9X=90
X=90/9=10